OpenStudy (anonymous):

The difference of two numbers is 6. The sum of the numbers is 28. Find the numbers. I need to see this worked out please.

OpenStudy (anonymous):

simulatous equation, a-b = 6 a+b= 28 add each part to the other 2a = 34 a=17 17- b = 6 -b= -11 b = 11

OpenStudy (sburchette):

First, we need to assign a variable to each unknown. We can say that x is one number and y is the other. The problem tells us that the difference of the two numbers is 6. That means x-y=6. It also says that their sum is 28, so x+y=28. We can use elimination to solve the equations. If we add the equations together, we have that x+x-y+y=6+28. So, 2x=34. So, x=17. We can substitute x=17 in either equation. If we use x-y=6, we have 17-y=6. So, y=11. The two numbers, then, are 17 and 11.

OpenStudy (anonymous):

Thank you SBurchette!! That makes more sense to me. I tried every way to figure it out thank you so much